OpenStudy (zab505):

The binomials shown below will be multiplied to produce a quadratic trinomial. (x + p)(x + q) Which part of the trinomial will equal the sum of p and q?

OpenStudy (zab505):

A. The constant term B. The coefficient of the x-term C. The degree D. The coefficient of the x2-term

OpenStudy (jdoe0001):

let's see if we can find it... let's try an example, let's set p = 2 and q= 3 so \(\large (x+2)(x+3)\implies \begin{array}{rrllll} x+2\\ \times x+3 \\\hline\\ x^2+2x\\ 3x+6\\ \hline\\ x^2{\color{blue}{ +2x+3x}}+6\\ \hline\\ x^2{\color{blue}{ +5x}}+6 \end{array}\) see which one is it now?

OpenStudy (anonymous):

Wouldn't that be: (x + p)(x + q) x * x = x^2 x * q = xq p * x = xp p * q = pq x^2 + xq + xp + pq?

OpenStudy (zab505):

I'm lost?

OpenStudy (jdoe0001):

well.. do you know what a "constant term" is? and a "coefficient" ?

OpenStudy (anonymous):

Yet it's supposed to be a trinomial?

OpenStudy (anonymous):

Constant = Variable. Coefficient = Before variable. Example = 2c, with c being constant and 2 being coefficient.

OpenStudy (zab505):

Oh ok

OpenStudy (anonymous):

Degree = 2c^1, 2, 3, etc.

OpenStudy (anonymous):

X2 would be x^2.

OpenStudy (zab505):

ah

OpenStudy (zab505):

wait so it would be d?

OpenStudy (anonymous):

But are you sure it's (x + p)(x + q)? I'm not sure how that will equal a trinomial.

OpenStudy (zab505):

yea that's what it says

OpenStudy (anonymous):

And it is more likely to be B, I think, but I'm a bit confused as well.

OpenStudy (zab505):

ah ok, yeah the question is confusing